OpenStudy (anonymous):

I have the following problem 0.3(4x+5)-1.7(x-4)=6.8 I have come up with x=3 is this correct

OpenStudy (anonymous):

0.3(4x+5)-1.7(x-4)=6.8 1.2x + 1.5 -1.7x + 6.8 = 6.8 1.2x - 1.7x + 1.5 + 6.8 = 6.8 - 0.5x = 6.8 - 8.3 - 0.5x = -1.5 x = -1.5/-0.5 x = 3 so yr ans is correct...